OpenStudy (anonymous):

What is the equation of the line that passes through (4, 2) and is parallel to 3x – 2y = –6?

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (skullpatrol):

Please let the student do the "simple algebra," how else will they learn?

OpenStudy (anonymous):

A.y=-3/2x+8 B.y=3/2x-4 C.y=1/2x D.y=2/3x-4

OpenStudy (anonymous):